A person has three times as many two dollar coin in his pocket as one dollar coins.The total value of these coins is $49. How many coins of each kind has he?
Answer by DrBeeee(684)

Let x = the number of $2 coins
Let y = the number of $1 coins
We can write two equations, one for the number of coins and the other for the total value of the coins. For the number of coins we have
(1) x = 3*y
For the value we have
(2) 2*x + 1*y = 49
Now put x of (1) into (2) and get
(3) 2*(3*y) + y = 49 or
(4) 7*y = 49 or
(5) y = 7
Then using (1) we get
(6) x = 3*7 or
(7) x = 21
Let's check using (2).
Is (2*21 + 1*7 = 49)?
Is (42 + 7 = 49)?
Is (49 = 49)? Yes
Answer: The person has 21 two dollar coins and 7 one dollar coins.